OpenStudy (kitkat1):

The sets of numbers 3, 4, 5 and 8, 15, 17 are Pythagorean triples. Use what you know about the Pythagorean Theorem and explain or show why they are Pythagorean triples. Be sure to show your work for each set of triples!

OpenStudy (johnweldon1993):

Hint, what is the pythagorean theorem?

OpenStudy (kitkat1):

a^2 + B^2=C^2

OpenStudy (johnweldon1993):

Correct! So theoretically, if these sets are indeed pythagorean triplets, if we plug them into this formula we will get an expression that is true Example, with the first set 3,4,5 if we plug those in \[\large 3^2 + 4^2 = 5^2\] is that true?

OpenStudy (kitkat1):

yes

OpenStudy (johnweldon1993):

Then it checks out...check the 2nd as well

OpenStudy (kitkat1):

8^2 +15^2 =17^2

OpenStudy (kitkat1):

So that how I check if its a triple or not?

OpenStudy (johnweldon1993):

Go through it...what is 8^2? add that to what 15^2 is...and see if that equals 17^2

OpenStudy (kitkat1):

O i got it thanks